# Wave Function Collapse, Lorentz Invariance, and the Third Postulate of Relativity

### v3

@@ III Interaction-Induced Wave Function Collapse
In Section II it was shown that stochastic collapse equations provide a very natural extension of the mathematical formalism of quantum theory and that they entail the measurement postulates as straightforward consequences of the fundamental equation. This approach eliminates the need to insert them into the logical structure of the theory at some very vaguely defined point. In this section it was shown that it is possible to formulate such an equation that respects conservation laws in individual measurement processes without introducing any new, ad hoc physical constants, and is also Lorentz invariant. The next section will consider whether this general approach can be considered to respect “serious” Lorentz invariance.
IV Relativity, Nonlocal Quantum Effects, and Spacetime Ontology
@@
The advent of quantum theory radically altered this situation. Einstein was the first to recognize this, but it was Bell’s demonstration of the reality of nonlocal effects that has really forced us to reconsider the ontology of spacetime. It is time to acknowledge that the assumption of local commutativity constitutes a third postulate of relativity, and that, like Einstein’s two original postulates it should be explained in terms of spacetime structure. The nonlocal correlations should, of course, be explained in a natural way. The stochastic collapse equation described in Section III shows how this can be done.
@@
Because the proposal described here is formulated in a nonrelativistic framework it must be considered as incomplete. But since the proposed modification is Lorentz invariant there do not appear to be any serious conceptual obstacles to a full relativistic account. Such an account would encompass quantum field theory, and might well offer solutions to some of the key problems with this more limited proposal. As argued in [16] there is reason to believe that such an extension could resolve the small discrepancies with energy conservation that are implied by equation 10. One might also hope that a relativistic account might provide a solution to the tails problem (mentioned in a footnote in section II). With the transition from distinguishable to indistinguishable particles the suppression of the very small amplitudes in the wave function tail below the level of vacuum fluctuations might point the way to a resolution. But these are just speculations that must be addressed in future work.

### v4

@@ IV Lorentz Invariance
In Section II it was shown that stochastic collapse equations provide a very natural extension of the mathematical formalism of quantum theory and that they entail the measurement postulates as straightforward consequences of the fundamental equation. This general approach eliminates the need to insert them into the logical structure of the theory at some very vaguely defined point. In Section III it was shown that the particular modification of the Schrödinger equation proposed in [16] insures the maintenance of conservation laws in individual measurement processes without introducing any new, ad hoc physical constants. In this section it was shown that the collapse operator is Lorentz invariant. The next section considers what implications this demonstration has for our understanding of the relationship between nonlocal quantum effects and relativity.
V Relativity, Nonlocal Quantum Effects, and Spacetime Ontology
@@
The advent of quantum theory radically altered this situation. Einstein was the first to recognize this, but it was Bell’s demonstration of the reality of nonlocal effects that has really forced us to reconsider the status of relativity and the implications for the ontology of spacetime. In classical theory Lorentz invariance followed from Einstein’s two postulates, but this is no longer true in quantum theory. As emphasized by Weinberg, the assumption of local commutativity is essential to maintain this defining characteristic of a relativistic theory. It is time to acknowledge that it constitutes a third postulate of relativity, and that, like Einstein’s two original postulates it should be explained in terms of spacetime structure. The nonlocal correlations should, of course, be explained in a natural way. The stochastic collapse equation described in Section III shows how this can be done.
@@
Because the proposal described here is formulated in a nonrelativistic framework it must be considered as incomplete. But since the proposed modification is Lorentz invariant there do not appear to be any serious conceptual obstacles to a full relativistic account. Such an account would encompass quantum field theory, and might well offer solutions to some of the key problems with this more limited proposal. As argued in [16] there is reason to believe that such an extension could resolve the small discrepancies with energy conservation that are implied by equation 10. One might also hope that a relativistic account might provide a solution to the tails problem (mentioned in a footnote in section II). With the transition from distinguishable to indistinguishable particles the suppression of the very small amplitudes in the wave function tail below the level of vacuum fluctuations might point the way to a resolution. These possibilities will be addressed in future work.